DCrandall13:

y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

Sunnyboi45:

Well if Y= x-1 then plug into Y in the problem y=3x+3 and solve it using common factors and terms.

jhonyy9:

given y = 3x+3 and y = x-1 but we know y = y so in this way we can write 3x+3=x-1 solve this for x